### v1

# LangTime: A Language-Guided Unified Model for Time Series Forecasting with Proximal Policy Optimization

Abstract
Recent research has shown an increasing interest in utilizing pre-trained large language models (LLMs) for a variety of time series applications. However, there are three main challenges when using LLMs as foundational models for time series forecasting: (1) Cross-domain generalization. (2) Cross-modality alignment. (3) Error accumulation in autoregressive frameworks. To address these challenges, we proposed LangTime, a language-guided unified model for time series forecasting that incorporates cross-domain pre-training with reinforcement learning-based fine-tuning. Specifically, LangTime constructs Temporal Comprehension Prompts (TCPs), which include dataset-wise and channel-wise instructions, to facilitate domain adaptation and condense time series into a single token, enabling LLMs to understand better and align temporal data. To improve autoregressive forecasting, we introduce TimePPO, a reinforcement learning-based fine-tuning algorithm. TimePPO mitigates error accumulation by leveraging a multidimensional rewards function tailored for time series and a repeat-based value estimation strategy. Extensive experiments demonstrate that LangTime achieves state-of-the-art cross-domain forecasting performance, while TimePPO fine-tuning effectively enhances the stability and accuracy of autoregressive forecasting.
Large language models, Time series, Reinforcement learning
1 Introduction
Time series refers to sequences of data points indexed in discrete-time order (Box et al., 2015), and they are common in real-world applications, such as financial risk assessment (Baffour et al., 2019), energy sustainability (Yu et al., 2023), and weather forecasting (Yu et al., 2025; Sun et al., 2021). Recently, large language models (LLMs) (Radford, 2018) have demonstrated remarkable capabilities in capturing sequential structures and patterns, which is crucial for modeling time-dependent data in time series forecasting. Given the similarities between time series and natural language in sequence modeling, several approaches have successfully leveraged pre-trained language models for time series forecasting, yielding promising outcomes (Xue & Salim, 2023; Gruver et al., 2024; Pan et al., 2024; Jia et al., 2024).
(a) Challenge 1: Cross-domain generalization.
(b) Challenge 2: Cross-modality alignment.
(c) Challenge 3: Error accumulation.
Figure 1: Challenges of Applying LLMs to Time Series Data. (a) Cross-domain generalization: Different domains possess their unique characteristics and numerical implications. (b) Cross-modality alignment: LLMs struggle to directly comprehend unseen time series as they are trained on language data, whereas language guidance can assist in enhancing their understanding. (c) Error accumulation: In autoregressive prediction, using the output from the previous step as input is considered unreliable.
However, as shown in Figure 1, arising from the inherent differences between time series and natural language, applying LLMs to time series forecasting presents three primary challenges. Firstly, cross-domain generalization. In contrast to natural language, which is governed by structural and semantic rules that ensure consistency across various domains, time series display a wide array of statistical characteristics across different fields. Furthermore, identical values may possess disparate meanings across distinct domains, thereby complicating the task of integrating multi-domain information into LLMs effectively. To address this challenge, some methods leverage semantic information embedded in datasets as prompts, enabling LLMs to effectively differentiate and adapt to various domains (Liu et al., 2024d). Secondly, cross-modality alignment. Pre-trained on an extensive text corpus, LLMs exhibit limited capacity for directly understanding time series, thereby rendering cross-modal alignment a formidable challenge (Zhou et al., 2023). While language as prefixes partially integrates temporal information, they fall short of achieving seamless alignment between the two modalities, limiting meaningful interaction and effective integration (Jin et al., 2023).
In addition, error accumulation in autoregressive frameworks. For tasks with varying prediction horizons, autoregressive models allow a single framework to handle multiple horizons, avoiding the need for separate training protocols (Liu et al., 2024f, e). However, supervised training is limited to optimizing a model’s ability to predict the next step based on actual historical observations. As a result, it fails to alleviate the adverse effects of error accumulation in autoregressive prediction.
To address these challenges, we propose LangTime, a language-guided unified model for time series forecasting that incorporates cross-domain pre-training and reinforcement learning-based fine-tuning. Specifically, we design Temporal Comprehension Prompts (TCPs) to integrate semantic information and channel details to help LLMs differentiate time series across various domains. Additionally, we condense the time series data into a single token and introduce a reconstruction task to enhance the understanding of the temporal patterns of LLM. Furthermore, we propose TimePPO, a fine-tuning algorithm based on Proximal Policy Optimization (PPO), which mitigates error accumulation during testing and improves long-term forecasting performance. Our contributions are summarized as follows:
  • We propose LangTime, an autoregressive model that integrates Temporal Comprehension Prompts to provide domain and channel-specific information, enabling LLMs to better understand and forecast time series.
  • We introduce TimePPO, a novel fine-tuning algorithm that alleviates error accumulation in autoregressive predictions, improving long-term forecasting accuracy.
  • Extensive experiments demonstrate that LangTime achieves state-of-the-art performance on widely recognized benchmarks and exhibits strong transferability to unseen domains. Our code are publicly available at: https://github.com/niuwz/LangTime.
2 Related Work
2.1 Large Language Models for Time Series
LLMs have shown considerable promise through pre-training on data spanning various domains (Doddapaneni et al., 2021; Taylor et al., 2022; Zhan et al., 2024). Nonetheless, foundational models for time series face substantial challenges in aligning multi-domain data due to differences in channel numbers, sampling frequencies, and patterns. Existing works address this issue using tailored strategies. TTM (Ekambaram et al., 2024) employs frequency prefixes, Lag-Llama (Rasul et al., 2023) uses lag features as covariates for probabilistic univariate forecasting, UniTime (Liu et al., 2024d) introduces masking and domain instructions to mitigate convergence imbalances, and ROSE (Wang et al., 2024) adopts frequency-based masking and reconstruction to unify cross-domain representations.
While multi-domain approaches focus on aligning time series data across sources, recent methods explore integrating textual information to further enhance forecasting capabilities, introducing a new challenge of cross-modality alignment. Directly concatenating time series and language tokens often lead to a modality gap due to structural and semantic differences. To address this, CALF (Liu et al., 2024c) applies knowledge distillation, S2IP-LLM (Pan et al., 2024) aligns semantic and time series spaces with tokenization and anchors, and TimeCMA (Liu et al., 2024a) uses prompt-based techniques to extract time series representations.
However, existing methods address these challenges in isolation, leaving a gap in simultaneously solving multi-domain and cross-modality alignment. LangTime bridges this gap by tackling both challenges concurrently. By facilitating robust cross-modality alignment, LangTime achieves superior performance in time series forecasting across diverse domains.
Figure 2: Overview of LangTime: (1) LangTime constructs Temporal Comprehension Prompts to help LLMs understand <Time Series Representation>, which is generated by the Temporal Encoder. (2) Alignment of LLMs with Time series during the pre-training via compressed representation learning and reconstruction-prediction tasks. (3) TimePPO fine-tune algorithm to mitigate error accumulation and enhance long-term forecasting. The content of the prompt has been simplified in the figure, with the detailed information provided in Section 3.3 and Section A.2.
2.2 Reinforcement Learning from Human Feedback in Large Language Models
Reinforcement Learning from Human Feedback (RLHF) has gained widespread application and demonstrated remarkable success in Natural Language Processing (NLP) tasks. For example, InstructGPT (Ouyang et al., 2022) uses human preference data to train a Reward Model (RM), which is subsequently employed to fine-tune the supervised policy using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017). This framework effectively enhances model performance by aligning outputs with human preferences. However, the requirement to train and optimize a reward model introduces significant computational overhead, making PPO resource-intensive in practical scenarios.
Despite the success of RLHF in NLP, its application to time series tasks remains underexplored, especially for autoregressive models. Time series forecasting introduces unique error accumulation challenges that require significant adaptation of existing RLHF methods. To bridge this gap, we adapt the PPO algorithm to the specific demands of time series prediction tasks. We redesign the reward function and value function to better capture the temporal structure and mitigate error propagation, simplifying the algorithm while enhancing its applicability to complex temporal data. These innovations make PPO more efficient and effective for autoregressive time series forecasting, addressing challenges such as computational complexity and scalability in long-term predictions.
3 Methodology
Problem Definition. Given the historical observations of multivariate time series $\mathbf{X}_{t}=\{\mathbf{x}_{t-L:t}^{i}\}_{i=1}^{C}$, where $L$ represents the number of lookback time steps and $C$ denotes the number of variates, the goal is to predict the future $F$ time steps $\mathbf{\hat{Y}}_{t}=\{\mathbf{\hat{x}}_{t:t+F}^{i}\}_{i=1}^{C}$. The ground truth of the future values is denoted as $\mathbf{Y}_{t}=\{\mathbf{x}_{t:t+F}^{i}\}_{i=1}^{C}$. LangTime is pre-trained on multi-source datasets $\mathcal{D}_{\text{pre-train}}=\{(\mathbf{X}_{t}^{j},\mathbf{Y}_{t}^{j})\}^{N}_{j=1}$, where $N$ is the number of datasets. For downstream task in dataset $j$, the model is fine-tuned on $\mathcal{D}^{j}_{\text{fine-tune}}$ and tested on $\mathcal{D}^{j}_{\text{test}}$. The datasets for pre-training, fine-tuning, and testing are pairwise disjoint to ensure generalization.
3.1 Architecture
LangTime includes three novel components: the Temporal Encoder (TE), Temporal Comprehension Prompts (TCPs), and Time Series Proximal Policy Optimization (TimePPO), as illustrated in Figure 2. To extract meaningful representations, LangTime employs the TE to process continuous time series. To bridge the gap between time series and language, TCPs encode essential contextual information, enabling LLMs to effectively differentiate and interpret time series data. The processed time series representations and contextual information are then passed into a pre-trained LLM. To further improve forecasting stability, LangTime incorporates TimePPO, a reinforcement learning-based fine-tuning strategy designed to mitigate error accumulation and enhance multi-step prediction. These components work together to ensure effective alignment between time series and language and enhance long-term forecasting performance.
Pre-training. To align time series representation with the word embedding space, we utilize two complementary pre-training tasks: reconstruction and prediction. The reconstruction task enhances the model’s understanding of time series, while the prediction task leverages the generative capabilities of LLMs to identify anticipatory dependencies for future forecasting (Cao et al., 2020). We adopt the Huber loss (Huber, 1992) to balance robustness and sensitivity. The overall pre-training loss is formulated as:
$\mathcal{L_{\text{pre-train}}}=\alpha\mathcal{L}_{\text{reconstruction}}+(1-\alpha)\mathcal{L}_{\text{prediction}},$ (1)
where $\alpha$ controls the trade-off between the reconstruction and prediction tasks. Details about the loss function can be found in Section A.1.
Input Processing and Adaptive Training. The input sequence $X\in\mathbb{R}^{L\times C}$ is first divided into non-overlapping patches of length $P$. For each patch $\mathcal{P}_{i}\in\mathbb{R}^{C\times P}$, we apply a linear transformation to project it into $\mathcal{P}^{\prime}_{i}\in\mathbb{R}^{C\times D}$.
To enable long-term forecasting in an autoregressive manner, our model maintains a fixed output sequence length while allowing the input sequence length to vary, constrained to integer multiples of the patch size. This design exposes the model to diverse input lengths, improving its generalization ability for autoregressive predictions. To address the imbalance caused by longer input sequences overshadowing shorter ones, we propose a progressive training strategy. This approach transitions from short to long sequence data, allowing the model to gradually adapt to the complexities of longer sequences.
3.2 Temporal Encoder
To align time series with LLMs, we introduce a channel-independent lightweight TE (Nie et al., 2022) to capture temporal dependencies and generate rich time series representations, which are mapped into the word embedding space using a simple linear layer (Liu et al., 2024b):
$F=\text{Linear}(\text{TE}(\mathcal{P}^{\prime})).$ (2)
Time series datasets exhibit varying convergence rates due to differences in their underlying statistical properties (Liu et al., 2024d). To address this, we introduce a random masking strategy to enhance the generalization across diverse datasets of LLMs. Specifically, we use a learnable token $\mathcal{P}_{m}$ to randomly replace patches in $F$ at a proportion of $r_{m}$, thereby obtaining $F_{m}$. During training, portions of the time series representation are randomly masked, prompting the model to infer missing values and capture deeper temporal structures for the reconstruction task. This strategy not only strengthens the model’s robustness but also improves its adaptability to datasets with distinct characteristics.
3.3 Temporal Comprehension Prompts
To fully leverage the generative capabilities of LLMs, we introduce TCPs specifically designed to bridge the gap between time series and language models. TCPs align time series representations with the structural characteristics of LLMs, facilitating both comprehension and forecasting. A detailed illustration of TCPs is provided in Section A.2.
Temporal Comprehension Prompts The information of the given time series:
Period: <Timestamp>,
Dataset: <Dataset Information>,
Channel: <Channel Information>,
Time: <Time Series Representation>,
Please compress this series into one word: <|EMB|>.
Based on the given information, predict next <N> values: <|OUT|>.
TCPs play a crucial role in guiding the LLM’s interpretation of time series data by encoding essential contextual information. The domain description encodes dataset-specific characteristics, allowing the model to incorporate relevant contextual information. The time placeholder is replaced with time series features extracted by TE to provide direct temporal input. Additionally, compression token enables the condensation of time series information into a single token, enhancing the model’s ability to capture key temporal patterns. The prediction guidance directs the forecasting process, ensuring smooth integration between time series data and the LLM’s generative structure.
All TCP components are fed into the pre-trained LLM, where the causal attention mechanism enables each token to incorporate information from all preceding tokens. Following (Zhang et al., 2024), we extract the previous token from <|EMB|>  and <|OUT|>  respectively, serving as the compressed token and prediction token. Ultimately, we employed a projection operation to obtain the reconstruction and prediction:
$\hat{X}$ $=\text{Linear}(\text{LLM}(F_{m})[\text{index}(\textless|\text{EMB}|\textgreater)-1]),$ (3)
$\hat{Y}$ $=\text{Linear}(\text{LLM}(F_{m})[\text{index}(\textless|\text{OUT}|\textgreater)-1]).$ (4)
The compressed token serves as a global summary of the time series and is used for reconstruction, while distinct linear layers handle variable sequence lengths, allowing the model to generalize across different history lengths. Under the guidance of TCPs, LLMs forecast future values by referencing both temporal embeddings and the compressed token, ensuring seamless integration of temporal and linguistic representations. This design enhances alignment between time series and LLMs, allowing LangTime to leverage LLMs’ strengths in sequential modeling while enhancing forecasting performance.
3.4 Time Series Proximal Policy Optimization
Figure 3: Overview of TimePPO. Both the Reference Model and the Fine-Tuned Model are initialized with the same parameters.
Autoregressive methods often suffer from high variance due to error propagation, which significantly degrades performance in long-term forecasting compared to non-autoregressive methods (Taieb & Atiya, 2015). To address this issue, we propose TimePPO, an extension of the PPO algorithm (Schulman et al., 2017) tailored for time series forecasting. Considering the unique characteristics of time series, we have undertaken a redesign of the Rewards Function and Value Function.
Rewards Function. As shown in Figure 3, the rewards function evaluates the discrepancy between the predicted series $\hat{y}_{t}$ and the ground truth $y_{t}$ at time $t$ from multiple perspectives, Additionally, we introduce an MSE penalty term $MSE(\hat{y}_{t},\hat{y}_{t}^{\text{ref}})$ to constrain the difference between the new policy and the old policy, ensuring stability and safety during policy updates, as defined in Equation 5:
$R(\hat{y}_{t})=$ $\text{tanh}\left(\tau\sum_{i}\mathcal{R}_{i}(\hat{y}_{t},y_{t})w_{i}\right)$ (5)
$-\beta MSE(\hat{y}_{t},\hat{y}_{t}^{\text{ref}}),$
where $\tau$ is a hyperparameter controlling the distribution of the reward scores, and $\mathcal{R}_{i}(\hat{y}_{t},y_{t})$ represents individual evaluation dimension (detailed in Section A.3). The weights $w_{i}$ balance the contributions of these dimensions.
$V(x_{t})=\begin{cases}0,&\text{if }t=0\\ \sum\limits_{i=t}^{m}\left(\text{tanh}\left(\tau\sum\limits_{j}\mathcal{R}_{i}(\hat{y}_{t-1},y_{i})w_{j}\right)\right).&\text{if }t>0\end{cases}$ (6)
Value Function. We define the value function $V(x_{t})$ to estimate the expected return from state $x_{t}$. As shown in Equation 6, we assume that the model will continue to repeat the output from time $t-1$ until the final time $m$, and then consider all the reward scores obtained as the expected return for predicting the entire sequence. This design ensures that the value function reflects the accumulated rewards, facilitating more accurate return estimation over time.
Advantage Estimation. To estimate the advantage, we employ Generalized Advantage Estimation (GAE) (Schulman et al., 2015):
$\hat{A}_{t}=\delta_{t}+(\gamma\lambda)\delta_{t+1}+\cdots+(\gamma\lambda)^{T-t+1}\delta_{T-1},$ (7)
$\delta_{t}=r_{t}+\gamma V\left(x_{t+1}\right)-\xi V\left(x_{t}\right),$
where $\xi$ is introduced to control the sensitivity to estimation errors in the current state’s value, thereby mitigating the error accumulation.
TimePPO. Inspired by InstructGPT, we introduce an alignment tax to prevent performance degradation (Ouyang et al., 2022). The loss function for TimePPO is given as:
$\mathcal{L}_{\text{TimePPO}}=$ $\,\mathbb{E}\left[\min\left(r(\theta)\hat{A}(x,y),\right.\right.$
$\left.\left.\text{clip}(r(\theta),1-\epsilon,1+\epsilon)\hat{A}(x,y)\right)\right]$
$+\eta\mathbb{E}||y-\hat{y}||_{2}^{2},$ (8)
where $r(\theta)$ represents the policy ratio, and $\epsilon$ regulates the extent of policy updates to ensure stability. The alignment tax term $\eta\mathbb{E}||y-\hat{y}||_{2}^{2}$ penalizes large deviations from the ground truth, ensuring that the model remains aligned with accurate predictions. Further details are presented in Section A.4.
With the aforementioned design, we have successfully enhanced the PPO algorithm for time series, thereby optimizing the autoregressive prediction model from an entirely novel perspective. The overall process of the TimePPO is depicted in Algorithm 1.
Algorithm 1 TimePPO
1:  Input: initial policy parameters $\theta_{0}$ and $\theta_{\text{ref}}$.
2:  for $k=0,1,2,...$ do
3:     Collect set of trajectories ${\mathcal{D}}_{k}=\{x_{t},\hat{y}_{t},\hat{y}^{\text{ref}}_{t}\}$ by running policy $\pi_{k}=\pi(\theta_{k})$ and $\pi_{\text{ref}}=\pi(\theta_{\text{ref}})$.
4:     Compute rewards $R(\hat{y}_{t})$ and values $V(x_{t-1})$.
5:     Compute advantage estimates $\hat{A}_{t}$.
6:     Update the policy $\pi_{k}=\pi(\theta_{k})$ by maximizing the TimePPO objective in Section 3.4.
7:  end for
Table 1: Forecasting performance comparisons. The input sequence length is set to 96. The predictive lengths are set to 96, 192, 336, 720. Avg is averaged over all predictive lengths. Red: best performance for the entire row. Bold: best performance among models trained across datasets.
Method Models Trained Across Datasets Models Trained/Fine-tuned on Each Dataset
[…]
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.323 0.346 0.350 0.385 0.914 0.590 0.319 0.348 0.337 0.374 0.530 0.519 0.359 0.381 0.327 0.361 0.327 0.363 0.329 0.367 0.338 0.375
192 0.372 0.376 0.384 0.401 0.966 0.616 0.368 0.375 0.373 0.393 0.664 0.583 0.383 0.393 0.368 0.381 0.368 0.383 0.367 0.385 0.374 0.387
336 0.419 0.403 0.420 0.423 0.935 0.612 0.413 0.402 0.405 0.415 0.836 0.652 0.416 0.414 0.401 0.405 0.400 0.405 0.399 0.410 0.410 0.411
720 0.491 0.443 0.473 0.455 0.954 0.630 0.487 0.439 0.465 0.448 1.297 0.782 0.483 0.449 0.465 0.436 0.461 0.439 0.454 0.439 0.478 0.450
Avg 0.401 0.392 0.407 0.416 0.942 0.612 0.397 0.391 0.395 0.408 0.832 0.634 0.410 0.409 0.390 0.396 0.389 0.398 0.387 0.400 0.400 0.406
ETTm2 96 0.184 0.258 0.183 0.264 0.269 0.331 0.188 0.258 0.179 0.264 0.722 0.588 0.193 0.280 0.177 0.262 0.178 0.264 0.175 0.259 0.187 0.267
192 0.245 0.300 0.246 0.307 0.326 0.364 0.245 0.297 0.246 0.307 0.898 0.673 0.257 0.318 0.245 0.305 0.246 0.307 0.241 0.302 0.249 0.309
336 0.308 0.339 0.310 0.346 0.379 0.393 0.301 0.336 0.308 0.346 1.207 0.792 0.317 0.353 0.304 0.342 0.309 0.349 0.305 0.343 0.321 0.351
720 0.410 0.399 0.413 0.405 0.473 0.442 0.402 0.393 0.410 0.406 1.708 0.952 0.419 0.411 0.400 0.397 0.410 0.408 0.402 0.400 0.408 0.403
Avg 0.287 0.324 0.288 0.331 0.362 0.383 0.284 0.321 0.286 0.331 1.134 0.751 0.297 0.341 0.282 0.327 0.286 0.332 0.281 0.326 0.291 0.333
ETTh1 96 0.394 0.395 0.525 0.500 0.417 0.408 0.391 0.388 0.480 0.473 0.381 0.401 0.398 0.410 0.423 0.425 0.385 0.402 0.414 0.419 0.384 0.402
192 0.439 0.420 0.544 0.511 0.484 0.444 0.429 0.419 0.505 0.487 0.435 0.434 0.451 0.440 0.464 0.446 0.432 0.425 0.460 0.445 0.436 0.429
336 0.464 0.442 0.576 0.529 0.529 0.468 0.462 0.440 0.536 0.504 0.480 0.459 0.508 0.471 0.499 0.461 0.467 0.447 0.501 0.466 0.491 0.469
720 0.462 0.449 0.577 0.548 0.549 0.494 0.458 0.445 0.531 0.520 0.499 0.478 0.483 0.478 0.505 0.487 0.472 0.466 0.500 0.488 0.521 0.500
Avg 0.440 0.427 0.556 0.522 0.495 0.454 0.435 0.423 0.513 0.496 0.449 0.443 0.460 0.450 0.473 0.455 0.439 0.435 0.469 0.455 0.458 0.450
ETTh2 96 0.301 0.334 0.308 0.357 0.301 0.333 0.299 0.336 0.306 0.356 0.318 0.352 0.295 0.346 0.306 0.354 0.303 0.354 0.302 0.348 0.340 0.374
192 0.380 0.389 0.390 0.405 0.410 0.398 0.374 0.382 0.388 0.404 0.401 0.404 0.386 0.399 0.377 0.398 0.386 0.404 0.388 0.400 0.402 0.414
336 0.412 0.419 0.424 0.438 0.420 0.419 0.410 0.418 0.423 0.436 0.446 0.441 0.447 0.443 0.423 0.435 0.430 0.438 0.426 0.433 0.452 0.452
720 0.422 0.439 0.435 0.454 0.439 0.444 0.418 0.426 0.433 0.453 0.460 0.459 0.428 0.444 0.431 0.447 0.433 0.452 0.431 0.446 0.462 0.468
Avg 0.379 0.395 0.389 0.414 0.393 0.399 0.375 0.391 0.388 0.412 0.406 0.414 0.389 0.408 0.384 0.409 0.388 0.412 0.387 0.407 0.414 0.427
Electricity 96 0.199 0.277 0.279 0.382 0.188 0.267 0.181 0.266 0.210 0.381 0.206 0.277 0.204 0.293 0.184 0.268 0.184 0.270 0.181 0.270 0.168 0.272
192 0.213 0.296 0.276 0.379 0.217 0.291 0.185 0.273 0.249 0.351 0.224 0.296 0.207 0.295 0.204 0.286 0.188 0.275 0.188 0.274 0.184 0.289
336 0.234 0.316 0.285 0.385 0.236 0.319 0.198 0.281 0.259 0.357 0.251 0.322 0.219 0.308 0.222 0.308 0.203 0.290 0.204 0.293 0.198 0.300
720 0.272 0.357 0.322 0.409 0.272 0.346 0.241 0.320 0.298 0.362 0.318 0.380 0.263 0.341 0.269 0.345 0.243 0.322 0.246 0.324 0.220 0.320
Avg 0.230 0.312 0.291 0.389 0.228 0.306 0.201 0.285 0.254 0.363 0.250 0.319 0.223 0.309 0.220 0.302 0.205 0.289 0.205 0.290 0.193 0.295
Exchange 96 0.086 0.205 0.124 0.254 0.133 0.253 0.089 0.201 0.118 0.246 0.087 0.202 0.083 0.203 0.087 0.208 0.084 0.201 0.088 0.205 0.107 0.234
192 0.175 0.300 0.218 0.338 0.253 0.357 0.175 0.298 0.212 0.332 0.178 0.298 0.178 0.299 0.178 0.302 0.178 0.299 0.176 0.299 0.226 0.344
336 0.329 0.412 0.367 0.443 0.390 0.452 0.329 0.409 0.360 0.438 0.329 0.413 0.328 0.415 0.338 0.422 0.343 0.422 0.301 0.397 0.367 0.448
720 0.854 0.696 0.913 0.728 0.931 0.730 0.852 0.690 0.904 0.723 0.792 0.675 0.854 0.696 0.819 0.681 0.803 0.671 0.901 0.714 0.964 0.746
Avg 0.361 0.403 0.406 0.441 0.427 0.448 0.361 0.400 0.399 0.435 0.347 0.397 0.361 0.403 0.356 0.403 0.352 0.398 0.367 0.404 0.416 0.443
Weather 96 0.184 0.203 0.181 0.222 0.219 0.245 0.178 0.202 0.182 0.220 0.188 0.227 0.195 0.233 0.180 0.221 0.183 0.223 0.177 0.218 0.172 0.220
192 0.216 0.250 0.226 0.261 0.298 0.310 0.211 0.245 0.226 0.263 0.234 0.266 0.240 0.269 0.229 0.261 0.230 0.262 0.218 0.259 0.219 0.261
336 0.275 0.293 0.280 0.299 0.337 0.338 0.269 0.286 0.279 0.323 0.288 0.305 0.293 0.306 0.285 0.301 0.285 0.302 0.278 0.297 0.280 0.306
720 0.361 0.349 0.356 0.347 0.415 0.383 0.351 0.346 0.354 0.346 0.363 0.355 0.368 0.354 0.359 0.349 0.362 0.351 0.354 0.348 0.365 0.359
Avg 0.259 0.274 0.261 0.282 0.317 0.319 0.252 0.270 0.260 0.288 0.268 0.288 0.274 0.291 0.263 0.283 0.265 0.285 0.257 0.281 0.259 0.287
$1^{\text{st}}$ Count 58 11 11 49 1 5 2 0 2 11 7
[…]
* indicates the adoption of the official baseline code, with adjustments to input sequence length and maximum training epochs for fair comparison with other methods, and other results are sourced from TimesNet (Wu et al., 2022).
4 Experiments
4.1 Training Details
Datasets. We extensively evaluate LangTime on seven real-world benchmark datasets spanning various time series application domains, i.e., ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, and Weather. The details of these datasets are shown in the Section B.1
Baselines. We evaluate LangTime against state-of-the-art models. (1) LLM-based methods: UniTime (Liu et al., 2024d), AutoTimes (Liu et al., 2024e), S2IP-LLM (Pan et al., 2024), Time-LLM (Jin et al., 2023), and GPT4TS (Zhou et al., 2023); (2) Specific models: PatchTST (Nie et al., 2022), and TimesNet (Wu et al., 2022). We adopt Qwen2-0.5B-Instruction (Yang et al., 2024) as our backbone. Detailed implementations are provided in Appendix B.
Setup. We adopt pre-training to fine-tuning dataset ratio of $\mathcal{D}_{\text{pre-train}}:\mathcal{D}_{\text{fine-tune}}=8:2$. To ensure a fair comparison, all methods maintained a fixed look-back window of $L=96$ and predicted future values with lengths of $F=\{96,192,336,720\}$.
Table 2: Comparisons of forecasting performance among various fine-tuning algorithms. The results are presented as averages over four forecasting horizons: 96, 192, 336, and 720. Bold: best results. Table 9 shows the full results.
Method ETTm1 ETTm2 ETTh1 ETTh2 Electricity Exchange Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
LangTime 0.401 0.392 0.287 0.324 0.440 0.427 0.379 0.395 0.230 0.312 0.361 0.403 0.259 0.274
[…]
AutoTimes 0.942 0.612 0.362 0.383 0.495 0.454 0.392 0.399 0.228 0.306 0.427 0.448 0.317 0.319
[…]
4.2 Main Results
Comparison with Other Forecasting Methods. Table 1 presents the overall forecasting performance of our model. The table is divided into two sections by vertical lines. On the left, models are pre-trained across multiple datasets, while on the right, models are either trained separately or fine-tuned for each dataset.
Pre-trained LangTime achieves remarkable improvements over baseline models that are also trained across datasets, achieving the best performance in 58 out of 70 entries. This highlights LangTime’s ability to generalize effectively across diverse time series distributions. On the right side of the table, results indicate that TimePPO fine-tuned LangTime achieves competitive performance, surpassing models trained individually on each dataset in 50 out of 70 entries, establishing new state-of-the-art results. These findings validate LangTime’s effectiveness in handling time series data with diverse characteristics. Furthermore, compared to other LLM-based forecasting methods, LangTime demonstrates significant advantages, reinforcing its capability to enhance the alignment between LLMs and time series data.
Comparison with Other Fine-tuning Methods. Table 2 presents a comparison between our proposed TimePPO algorithm and Supervised Fine-tuning (SFT). Across most datasets, models fine-tuned with TimePPO achieve superior predictive performance compared to those fine-tuned with SFT. This demonstrates the effectiveness of TimePPO in mitigating cumulative errors and enhancing the stability of autoregressive forecasting models.
Table 3: Ablation studies on various components of temporal comprehension prompts on ETTh1 and Weather datasets. Results are reported as averages over four forecasting horizons: 96, 192, 336, and 720. Full results are accessible in Table 10.
LG TS DI CI ETTh1 Weather
MSE MAE MSE MAE
0.436 0.436 0.267 0.284
0.440 0.438 0.269 0.288
0.442 0.440 0.272 0.292
0.442 0.443 0.274 0.293
Table 4: Ablation studies on various dimensions of Rewards Function on ETTh1 and Weather datasets. Results are reported as averages over four forecasting horizons: 96, 192, 336, and 720. Full results are accessible in Table 11.
Model ETTh1 Weather
MSE MAE MSE MAE
[…]
All dimensions 0.429 0.433 0.259 0.280
TimePPO w/o $\mathcal{R}_{\text{MSE}}$ 0.435 0.437 0.263 0.281
TimePPO w/o $\mathcal{R}_{\text{MAE}}$ 0.433 0.438 0.261 0.282
TimePPO w/o $\mathcal{R}_{\text{KL}}$ 0.435 0.438 0.262 0.282
4.3 Ablation Studies
Temporal Comprehension Prompts. We analyzed the impact of TCPs at different levels of detail by segmenting TCPs into the following components: ① Language Guidance (LG); ② Timestamp (TS); ③ Dataset Information (DI); ④ Channel Information (CI). The detailed results in Table 3 show that removing CI and DI leads to a decline in LangTime’s performance, while the impact of TS is relatively minor. Notably, even in the absence of supplementary metadata, using only LG still yields favorable results, further confirming the effectiveness of our proposed approach in aligning LLMs with time series data.
Rewards Function. We evaluate LangTime’s predictions across three different dimensions and compute the corresponding reward scores, with the results presented in Table 4. The findings indicate that incorporating all dimensions significantly enhances the performance of pre-trained LangTime. Conversely, removing any single dimension leads to a decline in TimePPO’s effectiveness, with Mean Squared Error (MSE) exerting the most significant impact compared to the other two dimensions.
Table 5: Comparison of Zero-Shot Performance of Pre-trained Models. The input sequence length is set to 96 for the Traffic dataset and 48 for the Illness to fit patch size 24. The predictive lengths are set to 24, 36, 48, 60 for Illness, and 96, 192, 336, 720 for Traffic. Bold: best results.
[…]
MSE MAE MSE MAE MSE MAE
Traffic 96 0.486 0.259 0.550 0.363 0.575 0.376
192 0.525 0.284 0.536 0.343 0.557 0.353
336 0.594 0.298 0.642 0.429 0.697 0.453
720 0.686 0.327 0.675 0.444 0.727 0.397
Avg 0.573 0.292 0.601 0.395 0.639 0.395
Illness 24 4.071 1.482 4.221 1.525 5.189 1.572
36 3.962 1.443 4.235 1.496 4.676 1.530
48 4.006 1.477 4.349 1.515 5.060 1.642
60 4.053 1.492 4.632 1.559 4.667 1.603
Avg 4.023 1.474 4.359 1.524 4.898 1.587
4.4 Zero-Shot Transferability Analysis
In this section, we conduct an in-depth analysis of the zero-shot performance of our pre-trained model compared to other cross-domain training models in unseen domains. Table 5 demonstrates that LangTime consistently outperforms baseline models in most cases. This highlights the effectiveness of TCPs in enhancing the model’s adaptability to diverse, unseen time series distributions.
4.5 Parameter Analysis
The Sensitivity of $\alpha$. Figure 4 presents the impact of $\alpha$ on model performance. When only the prediction task ($\alpha=0$) or only the reconstruction task ($\alpha=1$) is applied, the model exhibits suboptimal performance. However, when balancing both tasks ($\alpha=0.4$ or $\alpha=0.6$), LangTime achieves the best results across both datasets. This highlights the importance of the reconstruction task in guiding the LLM’s understanding of time series structures.
(a) ETTh1
(b) Weather
Figure 4: Parameter sensitivity analysis on $\alpha$ used in Equation 1.
(a) ETTh1
(b) Weather
Figure 5: Parameter sensitivity analysis on $\tau$ used in Equation 5.
The Sensitivity of $\tau$. Figure 5 illustrates the impact of $\tau$ on TimePPO’s rewards function. Since $\tau$ influences the distribution of reward scores, its effect varies based on the dataset’s MSE and MAE convergence values. For ETTh1, performance remains relatively stable across different values of $\tau$, indicating low sensitivity. In contrast, for Weather, an improper $\tau$ selection leads to noticeable fluctuations, occasionally causing performance degradation after fine-tuning. Nevertheless, in most cases, $\tau$ enhances performance, demonstrating the robustness of our proposed rewards function.
4.6 Model Analysis
T-SNE Visualization. To mitigate channel confusion in training, we incorporate channel-wise instructions in TCPs. To further analyze its impact, we compare the mean pooling of temporal representations and compressed tokens generated by LLMs for the first three channels of the ETTh1 and Weather datasets using T-SNE visualization.
As shown in Figure 6, the TE processes all channels uniformly, resulting in overlapping representations across domains, particularly between High Useful Load and High Useless Load in ETTh1. However, with the guidance of TCPs, LLMs successfully differentiate these channels and capture their semantic relationships. For instance, Air Temperature and Potential Temperature exhibit stronger correlations compared to Air Pressure, reflecting intrinsic dependencies within the dataset.
(a) Temporal Representations
(b) Compressed Token
Figure 6: T-SNE visualization of the temporal representations and compressed token.
Figure 7: Attention map visualization for the last token. Special tokens and system prompts are hidden for a more intuitive display.
Attention Map Analysis. To further examine how LLMs interpret and integrate time series representations, we visualize the last token attention weights from the final layer of the LLM in LangTime, as shown in Figure 7. The results indicate that: ① The last token focuses primarily on the time series representations and the token preceding the compressed token placeholder, which plays a key role in reconstructing the input sequence. ② LLMs also attend to contextual and prediction-relevant information, reinforcing the effectiveness of TCPs in aligning time series with LLMs. These findings confirm that TCPs enhance LLMs’ ability to differentiate time series patterns, preserve domain-specific knowledge, and capture meaningful temporal dependencies.
5 Conclusion
In this work, we propose LangTime, a novel generalized model designed to address the challenges of leveraging multi-domain datasets for improving downstream time series forecasting. By incorporating TCPs, LangTime enhances the LLM’s ability to interpret time series embeddings and generate accurate predictions. Additionally, we introduce TimePPO, a fine-tuning algorithm specifically designed for time series to effectively mitigate error accumulation in long-term forecasting. Extensive experiments demonstrate that LangTime achieves state-of-the-art performance on standard benchmarks and exhibits strong zero-shot transferability to unseen domains. Future work will concentrate on broadening LangTime’s applicability to encompass a wider array of time series analysis tasks.
Appendix A Method Design Details
A.1 Pre-Training Loss Details
During pre-training, we combine two tasks, as shown in Equation 1. For each task, we use the Huber loss, as illustrated in Equation 9 and Equation 10.
$\mathcal{L}_{\text{reconstruction}}=\begin{cases}\frac{1}{2}(x-\hat{x},&\text{if }|x-\hat{x}|\leq\delta,\\ \delta\cdot(|x-\hat{x}|-\frac{1}{2}\delta),&\text{otherwise.}\end{cases}$ (9)
$\mathcal{L}_{\text{prediction}}=\begin{cases}\frac{1}{2}(y-\hat{y},&\text{if }|y-\hat{y}|\leq\delta,\\ \delta\cdot(|y-\hat{y}|-\frac{1}{2}\delta),&\text{otherwise.}\end{cases}$ (10)
A.2 Temporal Comprehension Prompts Details
Table 6: Dataset Information and Channel Information used in TCPs. i means column index in dataset Traffic and Electricity. The Traffic and Illness datasets are utilized to evaluate Zero-Shot performance, whereas the remaining datasets are employed for both training and testing.
Dataset Description Channel Description
ETT An hourly-sampled (minutely-sampled in ETTm1 and ETTm2) electricity transformer dataset intended for electrical asset monitoring, collected from one area in a province in China. High Useful Load
High Useless Load
Middle Useful Load
Middle Useless Load
Low Useful Load
Low Useless Load
Oil Temperature
Weather Meteorological indicator data with ten minute sample rate. Air Pressure
Air Temperature
Potential Temperature
Dew Point Temperature
Relative Humidity
Saturation Water Vapor Pressure
Actual Water Vapor Pressure
Water Vapor Pressure Deficit
Specific Humidity
Water Vapor Concentration
Air Density
Wind Velocity
Maximum Wind Velocity
Wind Direction
Precipitation
Duration of Precipitation
Short Wave Downward Radiation
Photosynthetically Active Radiation
Maximum Photosynthetically Active Radiation
Internal Logger Temperature
CO2 Concentration of Ambient Air
Exchange Daily exchange rates of the US dollar to eight different currencies ranging from 1990 to 2016. Exchange rate of the US dollar to the Australian dollar
Exchange rate of the US dollar to the British pound
Exchange rate of the US dollar to the Canadian dollar
Exchange rate of the US dollar to the Swiss franc
Exchange rate of the US dollar to the Chinese yuan
Exchange rate of the US dollar to the Japanese yen
Exchange rate of the US dollar to the New Zealand dollar
Exchange rate of the US dollar to the Singapore dollar
Electricity Hourly electricity consumption of 321 customers from 2012 to 2014. Electricity consumption of customer i
Traffic Hourly road occupancy rates data from 862 detectors across the freeways of the San Francisco Bay area, covering the years 2015 to 2016. Road occupancy rates detected by detector i
Illness Weekly influenza-like illness (ILI) patient data from US Centers for Disease Control (2002-2021) showing ILI patient ratio. Weight ILI Rate
Unweight ILI Rate
Number of ILI patients aged between 0-4 years old
Number of ILI patients aged between 5-24 years old
Total number of ILI patients across all age groups
Number of sentinel providers
Total patients
Table 6 provides a description of each channel in the datasets used in TCPs. The full prompts used in TCPs are as follows.
Temporal Comprehension Prompts <|im_start|>system
You are a helpful assistant, and your target is to summarize a time series and predict the next time series. Note: Value means the actual values of the time series, where each token represents data for <PATCH SIZE> consecutive time points. <|im_end|>
<|im_start|>user
The information of the given time series:
Period: <Timestamp>,
Dataset: <Dataset Information>,
Channel: <Channel Information>,
Time: <Time Series Representation>,
Please compress this series into one word: <|EMB|>.
Based on the given information, predict next <N> values: <|im_end|>
<|im_start|>assistant
<|OUT|><|im_end|>
In TCPs, <Timestamp> , <Dataset Information> , and <Channel Information>  define the domain description, encoding dataset-specific characteristics to help the model incorporate relevant contextual information. <Time Series Representation> , generated by the Temporal Encoder, represents the processed time series features. <|EMB|>  serves as a placeholder for the temporal series embeddings. <N>  specifies the number of patches predicted concurrently and <|OUT|>  represents the placeholder for the predicted outputs.
A.3 Rewards Dimension Details
Table 7: Dimensions of Rewards Function
Dimension Calculation Formula
MSE $\mathcal{R}_{\text{MSE}}=\frac{1}{\frac{1}{n}\sum_{t=1}^{n}(\hat{y}_{t}-y_{t})^{2}}$
MAE $\mathcal{R}_{\text{MAE}}=\frac{1}{\frac{1}{n}\sum_{t=1}^{n}|\hat{y}_{t}-y_{t}|}$
KL Divergence $\mathcal{R}_{\text{KL}}=\frac{1}{D_{\text{KL}}(P\parallel Q)}$
The Rewards Function we propose assesses the significance of the output across three distinct dimensions. As delineated in Table 7, we independently compute the Mean Squared Error (MSE), Mean Absolute Error (MAE), and KL divergence between the predicted outcomes and the ground truth. For KL divergence, we first compute the mean and variance based on the two sequences, and then calculate the KL divergence between the two continuous distributions. Subsequently, their reciprocals are employed as evaluation metrics. Ultimately, the reward score for the predicted outcomes is determined via Equation 5. The proximity of the predicted result $\hat{y}_{t}$ to the ground truth $y_{t}$ is directly proportional to the reward score attained.
A.4 TimePPO Details
Unlike the discrete action space in NLP, the output of time series prediction models is a sequence in a continuous space. Therefore, we have redesigned the loss function based on NLP to measure the prediction results of the model from the probability distribution in the continuous space corresponding to the time series.
TimePPO is specifically designed to handle the continuous action space of time series forecasting, where the ratio of predicted mean probabilities is computed. This design prevents instability arising from excessive policy updates, thereby enhancing robustness in long-term forecasting.
We consider time series prediction as a continuous action space and calculate the ratio of predicted mean probabilities:
$\pi(a|y)=\frac{1}{\sqrt{2\pi\sigma^{2}(y)}}\exp\left(-\frac{(a-\mu(y))^{2}}{2\sigma^{2}(y)}\right),$ (11)
where $\mu(y)$ represents the mean of sequence $y$, while $\sigma^{2}(y)$ denotes the variance of sequence $y$. Based on Equation 11, we employ the probability policy ratio $r(\theta)$ to quantify the variations between the two policies.
$r(\theta)=\frac{\pi_{\theta}(\mu\left(y_{\text{new}})|y_{\text{new}}\right)}{\pi_{\theta_{\text{old}}}\left(\mu(y_{\text{new}})|y_{\text{old}}\right)}.$ (12)
Based on this design, $r(\theta)$ can accurately reflect the impact on the distribution of prediction results before and after the model parameter updates. Compared with directly calculating the numerical error of prediction results, it exhibits stronger robustness.
Appendix B Experimental Details
B.1 Datasets
We perform comprehensive experiments on seven extensively employed time series datasets for long-term forecasting. Adhering to the data processing and train-validation-test set split protocol established in TimesNet (Wu et al., 2022), we ensure that the train, validation, and test datasets are rigorously partitioned in chronological order to prevent data leakage. Regarding the long-term forecasting configurations, we set the context length of LangTime and the lookback length of other comparative methods to 96, while the forecast length varies among {96, 192, 336, 720}. Brief descriptions of the pre-training datasets are as follows: (1) ETT (Zhou et al., 2021) includes data for monitoring electricity transformers from July 2016 to July 2018, comprising four subsets: ETTm1, ETTm2, ETTh1, and ETTh2. (2) Electricity contains hourly power consumption data for 321 clients from 2012 to 2014. (3) Exchange (Lai et al., 2018) records daily exchange rates for eight countries from 1990 to 2016. (4) Weather is recorded every 10 minutes in 2020, featuring 21 meteorological indicators such as temperature, humidity, and precipitation. Additionally, we evaluated the zero-shot performance of LangTime using two datasets from different domains: (1) Illness includes weekly data on patients with seven influenza-like illnesses from 2002 to 2021. (2) Traffic includes data on hourly road occupancy rates, gathered by 862 detectors across the freeways of the San Francisco Bay area from 2015 to 2016. Owing to the constraints of computational resources and given our adoption of a channel-independent approach, we partition the number of channels in certain datasets. Table 8 illustrates the number of channels present in each sub-dataset. This strategy facilitates more efficient management of batch size during cross-domain training.
Table 8: Summary of datasets used for pre-training.
Dataset Name Channels Frequency Time Points Channel Split Application Domain
ETTm1/ETTm2 7 15 mins 57,507 - Energy Infrastructure Observation
ETTh1/ETTh2 7 1 hour 14,307 - Energy Infrastructure Observation
Electricity 321 1 hour 26,211 7 Electricity Consumption
Weather 21 10 mins 52,603 7 Meteorologic Monitoring
Exchange 8 1 day 7,207 - Foreign Exchange Market
B.2 Experimental Setting
LangTime was implemented using PyTorch (Paszke et al., 2019), and all experiments were executed on 8 NVIDIA A100 80GB GPUs. We employed the AdamW optimizer (Loshchilov, 2017) with an initial learning rate of $1\times 10^{-4}$. A cosine annealing schedule with warmup was utilized for learning rate decay during pre-training, with a warmup rate set at 0.05. The Temporal Encoder in LangTime comprises 4 layers utilizing group query attention (Ainslie et al., 2023), with a query head count of 8, and key and value head counts of 2. The dimension of the latent space is set to 512. The patching process employs a patch size of 24, and LangTime predicts 4 patches simultaneously.
Pre-training. During the warmup phase, we established $\alpha=0.7$ to facilitate LLMs in comprehending the time series, subsequently reducing $\alpha$ to 0.5 during the learning rate decay phase. The Mask Rate was set at 0.4, and for the pre-trained model with a lookback length of 96, we designated the input lengths as {96, 288, 480, 672}. The batch size is set to 48 in pre-training.
Fine-tuning. Customizing the fine-tuning parameters can yield optimal results for different datasets. For instance, in the Weather dataset, we configure the initial learning rate to $1\times 10^{-6}$, maintain the lookback window at 96, and execute predictions with target lengths of 672, with a clip range $\epsilon$ set to 0.1. The parameter $\tau$ is set to 0.1, while $\xi$ is set to 0.9.
Appendix C Supplementary Results
C.1 Full Results
In this section, we present the experimental results that were not fully displayed in the main text. Specifically, Table 9 provides a detailed account of the impact of different fine-tuning algorithms on two models pre-trained across multiple domains. Table 10 and Table 11 display the complete results of the ablation experiments we conducted. Table 12 and Table 13 respectively display the experimental results for $\alpha$ and $\tau$ in the parameter sensitivity analysis.
Table 9: Comparisons of forecasting performance among various fine-tuning algorithms.
Method ETTm1 ETTm2 ETTh1 ETTh2 Electricity Exchange Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
[…]
192 0.372 0.376 0.245 0.300 0.439 0.420 0.380 0.389 0.213 0.296 0.175 0.300 0.216 0.250
336 0.419 0.403 0.308 0.339 0.464 0.442 0.412 0.419 0.234 0.316 0.329 0.412 0.275 0.293
720 0.491 0.443 0.410 0.399 0.462 0.449 0.422 0.439 0.272 0.357 0.854 0.696 0.361 0.349
Avg 0.401 0.392 0.287 0.324 0.440 0.427 0.379 0.395 0.230 0.312 0.361 0.403 0.259 0.274
[…]
192 0.366 0.374 0.245 0.296 0.447 0.418 0.384 0.390 0.198 0.286 0.183 0.305 0.221 0.253
336 0.417 0.403 0.304 0.337 0.478 0.437 0.416 0.423 0.210 0.293 0.325 0.412 0.284 0.299
720 0.488 0.439 0.406 0.395 0.461 0.448 0.424 0.429 0.249 0.323 0.856 0.695 0.377 0.356
Avg 0.399 0.391 0.285 0.321 0.447 0.423 0.378 0.391 0.211 0.291 0.362 0.404 0.263 0.279
[…]
192 0.368 0.375 0.245 0.297 0.429 0.419 0.374 0.382 0.185 0.273 0.175 0.298 0.211 0.245
336 0.413 0.402 0.301 0.336 0.462 0.440 0.410 0.418 0.198 0.281 0.329 0.409 0.269 0.286
720 0.487 0.439 0.402 0.393 0.458 0.445 0.418 0.426 0.241 0.320 0.852 0.690 0.351 0.346
Avg 0.397 0.391 0.284 0.321 0.435 0.423 0.375 0.391 0.201 0.285 0.361 0.400 0.252 0.270
[…]
192 0.966 0.616 0.326 0.364 0.484 0.444 0.410 0.398 0.217 0.291 0.253 0.357 0.298 0.310
336 0.935 0.612 0.379 0.393 0.529 0.468 0.420 0.419 0.236 0.319 0.390 0.452 0.337 0.338
720 0.954 0.630 0.473 0.442 0.549 0.494 0.439 0.444 0.272 0.346 0.931 0.730 0.415 0.383
Avg 0.942 0.612 0.362 0.383 0.495 0.454 0.392 0.399 0.228 0.306 0.427 0.448 0.317 0.319
[…]
192 0.956 0.605 0.326 0.365 0.480 0.438 0.432 0.416 0.217 0.284 0.253 0.357 0.299 0.310
336 0.940 0.606 0.383 0.393 0.526 0.461 0.420 0.419 0.238 0.317 0.389 0.452 0.338 0.338
720 0.969 0.626 0.477 0.444 0.544 0.493 0.448 0.444 0.274 0.367 0.932 0.732 0.415 0.383
Avg 0.945 0.613 0.364 0.383 0.491 0.448 0.400 0.403 0.228 0.307 0.424 0.444 0.318 0.319
[…]
192 0.961 0.598 0.333 0.365 0.476 0.438 0.410 0.406 0.221 0.286 0.256 0.360 0.298 0.310
336 0.938 0.605 0.380 0.403 0.517 0.465 0.416 0.417 0.237 0.308 0.388 0.458 0.337 0.338
720 0.948 0.627 0.471 0.443 0.528 0.479 0.436 0.439 0.267 0.345 0.928 0.727 0.415 0.378
Avg 0.940 0.610 0.360 0.383 0.485 0.446 0.390 0.399 0.233 0.304 0.425 0.450 0.317 0.318
Table 10: Ablation studies on various components of temporal comprehension prompts on ETTh1 and Weather datasets.
Language Guidance Timestamp Dataset Information Channel Information Length ETTh1 Weather
MSE MAE MSE MAE
96 0.384 0.399 0.175 0.209
192 0.432 0.429 0.235 0.265
336 0.471 0.449 0.293 0.312
720 0.459 0.467 0.364 0.349
Avg 0.436 0.436 0.267 0.284
96 0.385 0.402 0.179 0.214
192 0.434 0.429 0.238 0.267
336 0.476 0.451 0.295 0.315
720 0.465 0.471 0.366 0.356
Avg 0.440 0.438 0.269 0.288
96 0.392 0.407 0.180 0.217
192 0.441 0.431 0.240 0.267
336 0.468 0.450 0.299 0.318
720 0.469 0.471 0.369 0.365
Avg 0.442 0.440 0.272 0.292
96 0.393 0.407 0.180 0.220
192 0.440 0.433 0.239 0.269
336 0.470 0.454 0.306 0.320
720 0.467 0.477 0.369 0.365
Avg. 0.442 0.443 0.274 0.293
Table 11: Ablation studies on various dimensions of Rewards Function on ETTh1 and Weather datasets.
ETTh1 Weather
MSE MAE MSE MAE
[…]
192 0.432 0.429 0.235 0.265
336 0.471 0.449 0.293 0.312
720 0.459 0.467 0.364 0.349
Avg 0.436 0.436 0.267 0.284
All dimensions 96 0.379 0.399 0.172 0.205
192 0.426 0.427 0.226 0.263
336 0.466 0.448 0.288 0.310
720 0.445 0.458 0.352 0.342
Avg 0.429 0.433 0.259 0.280
TimePPO w/o $\mathcal{R}_{\text{MSE}}$ 96 0.382 0.401 0.175 0.208
192 0.430 0.429 0.229 0.264
336 0.472 0.452 0.289 0.311
720 0.456 0.465 0.359 0.342
Avg 0.435 0.437 0.263 0.281
TimePPO w/o $\mathcal{R}_{\text{MAE}}$ 96 0.381 0.401 0.172 0.208
192 0.429 0.430 0.225 0.264
336 0.470 0.453 0.289 0.312
720 0.453 0.468 0.357 0.345
Avg 0.433 0.438 0.261 0.282
TimePPO w/o $\mathcal{R}_{\text{KL}}$ 96 0.382 0.401 0.176 0.207
192 0.430 0.430 0.228 0.264
336 0.472 0.454 0.289 0.310
720 0.456 0.468 0.357 0.346
Avg 0.435 0.438 0.262 0.282
Table 12: Full result of parameter sensitivity analysis on $\alpha$
$\alpha$ Length ETTh1 Weather $\alpha$ Length ETTh1 Weather
MSE MAE MSE MAE MSE MAE MSE MAE
0 96 0.389 0.399 0.165 0.204 0.2 96 0.388 0.395 0.182 0.225
192 0.443 0.432 0.254 0.279 192 0.442 0.429 0.234 0.267
336 0.494 0.456 0.329 0.330 336 0.491 0.452 0.284 0.302
720 0.481 0.467 0.484 0.409 720 0.480 0.466 0.355 0.348
Avg 0.452 0.438 0.308 0.306 Avg 0.450 0.435 0.264 0.285
0.4 96 0.394 0.400 0.176 0.220 0.6 96 0.399 0.402 0.185 0.229
192 0.445 0.431 0.230 0.264 192 0.449 0.432 0.244 0.276
336 0.497 0.454 0.281 0.300 336 0.503 0.454 0.293 0.309
720 0.485 0.467 0.354 0.347 720 0.499 0.468 0.364 0.357
Avg 0.455 0.438 0.260 0.283 Avg 0.462 0.439 0.271 0.293
0.8 96 0.406 0.406 0.192 0.235 1 96 0.768 0.583 0.224 0.276
192 0.459 0.436 0.244 0.277 192 0.757 0.587 0.273 0.311
336 0.511 0.456 0.292 0.309 336 0.754 0.595 0.315 0.336
720 0.497 0.463 0.363 0.355 720 0.730 0.603 0.379 0.376
Avg 0.468 0.440 0.273 0.294 Avg 0.752 0.592 0.298 0.325
Table 13: Full result of parameter sensitivity analysis on $\tau$
$\tau$ Length ETTh1 Weather $\tau$ Langth ETTh1 Weather
MSE MAE MSE MAE MSE MAE MSE MAE
[…]
192 0.432 0.429 0.235 0.265 192 0.424 0.426 0.224 0.259
336 0.471 0.449 0.293 0.312 336 0.465 0.447 0.278 0.299
720 0.459 0.467 0.364 0.349 720 0.443 0.456 0.354 0.349
Avg 0.436 0.436 0.267 0.284 Avg 0.427 0.432 0.256 0.280
0.05 96 0.378 0.397 0.224 0.259 0.1 96 0.379 0.399 0.171 0.214
192 0.424 0.426 0.224 0.259 192 0.426 0.427 0.224 0.262
336 0.466 0.448 0.278 0.299 336 0.466 0.448 0.278 0.302
720 0.445 0.458 0.354 0.349 720 0.445 0.458 0.354 0.352
Avg 0.428 0.432 0.270 0.292 Avg 0.429 0.433 0.257 0.282
0.3 96 0.378 0.398 0.173 0.220 0.5 96 0.378 0.398 0.175 0.219
192 0.425 0.427 0.221 0.262 192 0.425 0.427 0.223 0.262
336 0.467 0.449 0.271 0.298 336 0.467 0.449 0.275 0.300
720 0.448 0.460 0.348 0.348 720 0.449 0.461 0.352 0.350
Avg 0.430 0.433 0.253 0.282 Avg 0.430 0.433 0.256 0.283
C.2 Cases
In this part, we visualize the forecasting results of LangTime. Figure 8 illustrates the performance comparison between pre-trained LangTime and LangTime fine-tuned via TimePPO under various settings. Figure 9 visualizes the zero-shot results of the pre-trained LangTime on the Traffic dataset. This demonstrates the exceptional performance of our proposed LangTime across domains, achieving impressive results even in unseen domains.
(a) Input-96-predict-96 results on ETTh1 dataset.
(b) Input-96-predict-192 results on ETTh1 dataset.
(c) Input-96-predict-96 results on Weather dataset.
(d) Input-96-predict-192 results on Weather dataset.
Figure 8: Long-term forecasting cases for ETTh1 and Weather dataset. Green lines are the ground truths, orange lines are the pre-trained model predictions and blue lines are the TimePPO fine-tuned model predictions. The vertical line indicates where the prediction starts.
(a) Input-96-predict-96 results on Traffic dataset.
(b) Input-96-predict-192 results on Traffic dataset.
(c) Input-96-predict-96 results on Traffic dataset.
(d) Input-96-predict-192 results on Traffic dataset.
Figure 9: Zero-Shot forecasting cases for Traffic dataset. Green lines are the ground truths, orange lines are the pre-trained model predictions. The vertical line indicates where the prediction starts.